5. Find ALL exact solutions in the interval [0,2 ) for the equation: v2 cos(2x) = -1
Answer #1

briven Equation Focos (29) = -1 COS(2x) = COS (23) = Cod (11- ) Cos (22) = (as (21) (do Cax then o anh ta if I nf I Here X =